What is the maximum guage pressure of water that can be measure by means of manometer tube, 2m high, density of water is 1000kg
1
Expert's answer
2021-07-15T13:16:16-0400

Gives

Hight(h)=2m

"\rho=1000kg/m^3"

We know that

Gauge pressure


"P=\rho gh=1000\times9.8\times2=19600pa"

Put value

"P_G=P-P_0\\P_G=\rho gh-P_0"

"P_0=1atm=10^5pa"

"P=19600-10^5=80.400pa\\P_G=80.400pa"